Tangents are drawn from a point M(3,6) to the circle    x2+y2=9. If point of contacts are P(a,b) and Q(c,d) respectively then find 15(a+b+c+d).

Grade:12

1 Answers

rusty
7 years ago
The slope of tangent at tangent point(x1,y1) would be -x1/y1. Then find the equation of tangent at point (x1,y1) passing through the given point (3,6).As the point (x1,y1) will also lie on circle, substitute the equation we will get in the equation of circle.By doing so, we will get the values of a=3, b= -9/5 , c=12/5, d=0Thus 15(a+b+c+d) = 54